# Between Lines of Code: Unraveling the Distinct Patterns of Machine and Human Programmers

### 12 Jan 2024 v1

@@ 5.6. Ablation Study (RQ3)
In our ablation study, we compared the effectiveness of different perturbation strategies for detecting machine-generated code using the CodeLlama (7B) model on The Stack dataset. The results summarized in Table 5 illuminates the comparative advantage of our structured perturbation approach. We observe that both newline and space perturbations independently offer substantial improvements over the traditional MLM-based (CodeT5+) perturbation technique as in DetectGPT and DetectLLM (Mitchell et al., 2023; Su et al., 2023) for natural language. Also, the combination of newline and space perturbations further enhances the detection performance, with the highest AUROC score of 0.8852 at $T=0.2$ and 0.6660 at $T=1.0$. The consistent outperformance of our combined perturbation strategy across both temperature settings affirms its potential as a robust solution for detecting machine-authored code.
5.7. Impact of Perturbation Count (RQ4)
@@
Table 6. Impact of varying the number of perturbations
@@
Results in Table 6 reveal a rapid ascent in the AUROC score as the number of perturbations increases from 10 to 20, underscoring the efficiency of our perturbation approach. Notably, an increase to 20 perturbations already yields robust detection performance, with further increments leading to diminishing improvements. This suggests that our method requires a relatively small number of perturbations to effectively discern between human- and machine-authored code. This implies that our method is not only effective but also resource-efficient.
5.8. Performance of Cross-Model Detection (RQ5)
Figure 5. Cross-model detection performance
@@ 5.8. Performance of Cross-Model Detection (RQ5)
The results presented in Figure 5 highlight DetectCodeGPT’s adaptability in cross-model detection. While the algorithm excels in the white-box setting, its performance endures with only a slight reduction when subjected to cross-model application. For instance, StarCoder, when detecting code generated by WizardCoder and CodeLlama, yields AUROC scores of 0.92 and 0.87, respectively, compared to an AUROC of 0.93 when detecting its own output. And we can also notice a performance decrease when detecting CodeGen2’s output. This is possibly due to the fact that CodeGen2 is trained on a more diverse dataset containing more natural language text (Nijkamp et al., 2023). However, Phi-1 demonstrates a relative proficiency with a score of 0.83 to detect CodeGen2’s output, which implies an ensemble of diverse detection models may enhance the system’s robustness, as suggested in (Mireshghallah et al., 2023).
These results indicate that DetectCodeGPT is a model-free method that is robust against model discrepancies.
6. Discussion
In this section, we will analyze why our method is effective and discuss its strengths and limitations.
6.1. Why is DetectCodeGPT Effective?
@@
Preservation of Code Correctness
@@
Emulation of Human Randomness
As discussed in Section 3.4.5, human inherently exhibit less naturalness and more randomness in their use of spaces and newlines than machines during coding. For example, a human programmer may freely insert whitespace, especially newlines, in the code as they deem fit, whereas a machine programmer usually tries to structure the code in a more standardized and compact manner. Our proposed perturbation strategy mimics human’s free usage of spaces and newlines, thereby making the perturbation more “random” as is desired according to Section 3.
6.2. Strength of DetectCodeGPT
DetectCodeGPT presents a perturbation strategy that is not only efficient but also maintains the structural integrity of the code while closely mirroring human coding preferences. We summarize the strengths of DetectCodeGPT as follows:
Enhanced Efficiency
Since it’s necessary to perturb each text for multiple times to ensure the effectiveness of detection (Mitchell et al., 2023; Su et al., 2023), as verified in Section 5.7, perturbing code using MLM will lead to huge computational cost. On the contrary, our method will save much more time and bring more efficiency. This translates to a significant conservation of computational resources, thus expediting the detection process and enhancing scalability. Mathematically, if $T_{\text{DetectGPT}}$ and $T_{\text{Ours}}$ denote the computational times of DetectGPT and our method respectively, we know that $T_{\text{Ours}}\ll T_{\text{DetectGPT}}$.
Superior Zero-Shot Performance
DetectCodeGPT distinguishes itself with a zero-shot learning capability, enabling it to detect machine-generated code without the necessity for training on extensive datasets. This model-agnostic advantage means that it can be generalized across various code LLMs and can still remain robust even when applied to unseen models, as demonstrated in Section 5.8. In comparison to the supervised methods, which may suffer from data drift or the need for constant updates, our approach is not only more adaptable but also delivers better performance across various data sources. This makes DetectCodeGPT an agile and reliable tool for machine-generated code detection in a variety of contexts.
6.3. Limitations and Future Directions
@@
Looking ahead to future work, we plan to further improve the effectiveness of DetectCodeGPT when detecting machine-generated code at higher levels of generation randomness, such as $T$=$1.0$. We note from Table 4 that although DetectCodeGPT outperforms other baselines at $T$=$1.0$, there is still large room for improvement. Although ensembling multiple detection models may help to improve the detection performance (Mireshghallah et al., 2023), we look forward to explore more effective perturbation strategies based on code structure to further enhance the detection efficacy.
@@
Additionally, the development of watermarking strategies (Sun et al., 2023) for machine-generated code represents an intriguing avenue for future research. Given that our method can effectively identify machine’s code, it could be adapted to assist in the creation of watermarking techniques. For instance, embedding a watermark could be achieved by inserting newlines and spaces in a code-specific pattern, which would increase detectablility without compromising functionality.
Moreover, our work also highlights the need for the development of more robust detection methods against possible attacks (Hu et al., 2023; Krishna et al., 2023). As discussed in Section 3, an important difference between machine- and human-authored code is naturalness. We can possibly leverage the insertion of whitespace to imitate a less ‘natural’ code style and thereby make machine’s code more difficult to detect. This potential strategy highlights the need for the development of more robust detection methods against possible attacks (Hu et al., 2023; Krishna et al., 2023).
7. Related Work
@@
Recently, there has been much effort in detecting machine-generated text (Yang et al., 2023; Wu et al., 2023b). The two main categories of detection methods are zero-shot and training-based methods, and our DetectCodeGPT falls into the former category.
@@
As for zero-shot methods, they are usually based on the discrepancy between likelihood and rank information of human and machine’s texts (Gehrmann et al., 2019; Ippolito et al., 2020; Solaiman et al., 2019). Leveraging the hypothesis in DetectGPT (Mitchell et al., 2023) that machine-generated text often has a negative curvature in the log probability when the text is perturbed, many perturbation based methods have been proposed (Mitchell et al., 2023; Su et al., 2023; Bao et al., 2023). These methods usually perturb the text by masking a continuous span of tokens and then recover the perturbed text using another LLM like T5 (Ahmad et al., 2021). The benefit of these methods is that they are zero-shot and can be applied to any LLM without access to training data. However, the perturbation process is time-consuming and computationally expensive,
When it comes to the training-based methods, fine-tuning the RoBERTa(Liu et al., 2019) or T5 (Ahmad et al., 2021) model with data collected from different model families at different decoding settings is a common practice (Tian et al., 2023; Zhan et al., 2023; Chen et al., 2023). Additional information like graph structure (Liu et al., 2022), perplexity from proxy models (Wu et al., 2023a) have been shown to be helpful for detection. Moreover, techniques like adversarial training (Hu et al., 2023) and contrastive learning (Liu et al., 2022) have also been proposed to improve the detection performance. The main challenge of training-based methods is that they often lack generalization ability and require access to training data from the target model.
@@
However, research on identifying machine-generated code remains relatively scarce and is purported to be more challenging than discerning machine-produced text (Lee et al., 2023). The predominant strategies in current studies largely revolve around watermarking techniques, which embed unique markers into the code either during the training of the model or at the time of code generation (Lee et al., 2023; Sun et al., 2023). The detection of these watermarks subsequently enables the recognition of code generated by machines. It should be noted, however, that these watermarking methods are primarily designed to address issues related to code licensing and plagiarism (Cox et al., 2002; Collberg and Thomborson, 1999). Their reliance on modifications to the generation model renders them unsuitable for general code detection tasks.
8. Conclusion
@@
In this paper, we have explored the nuanced differences between machine-generated and human-written code to enhance the trustworthiness of software development in the age of AI. By pioneering a comparative analysis across five different attributes, we have provided new insights into code provenance that machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in structural tokens such as whitespaces. Our proposed detection method, DetectCodeGPT, capitalizes on these insights, specifically targeting the structural nuances of code with a novel perturbation strategy that is simple yet effective. The empirical success of DetectCodeGPT against several benchmarks confirms its potential to bolster security measures in programming and establish a more symbiotic relationship between human developers and LLMs. Our work marks a significant step towards responsible AI use in software engineering, ensuring clarity in code authorship and maintaining the integrity of code as a craft.

### 24 Jan 2024 v2

@@ 5.6 Ablation Study (RQ2)
In our ablation study, we compared the effectiveness of different perturbation strategies for detecting machine-generated code using the CodeLlama (7B) model on The Stack dataset. The results summarized in Table V illuminates the comparative advantage of our stylistic perturbation approach. We observe that both newline and space perturbations independently offer substantial improvements over the traditional MLM-based (CodeT5+) perturbation technique as in DetectGPT and DetectLLM [11, 45] for natural language. Also, the combination of newline and space perturbations further enhances the detection performance, with the highest AUROC score of 0.8852 at $T=0.2$ and 0.6660 at $T=1.0$. The consistent outperformance of our combined perturbation strategy across both temperature settings affirms its potential as a robust solution for detecting machine-authored code.
5.7 Impact of Perturbation Count (RQ3)
@@
Table VI: Impact of varying the number of perturbations
@@
Results in Table VI reveal a rapid ascent in the AUROC score as the number of perturbations increases from 10 to 20, underscoring the efficiency of our perturbation approach. Notably, an increase to 20 perturbations already yields robust detection performance, with further increments leading to diminishing improvements. This suggests that our method requires a relatively small number of perturbations to effectively discern between human- and machine-authored code. This implies that our method is not only effective but also resource-efficient.
5.8 Performance of Cross-Model Detection (RQ4)
Figure 6: Cross-model detection performance
@@ 5.8 Performance of Cross-Model Detection (RQ4)
The results presented in Figure 6 highlight DetectCodeGPT’s adaptability in cross-model detection. While the algorithm excels in the white-box setting, its performance endures with only a slight reduction when subjected to cross-model application. For instance, StarCoder, when detecting code generated by WizardCoder and CodeLlama, yields AUROC scores of 0.92 and 0.87, respectively, compared to an AUROC of 0.93 when detecting its own output. And we can also notice a performance decrease when detecting CodeGen2’s output. This is possibly due to the fact that CodeGen2 is trained on a more diverse dataset containing more natural language text [4]. However, Phi-1 demonstrates a relative proficiency with a score of 0.83 to detect CodeGen2’s output, which implies an ensemble of diverse detection models may enhance the system’s robustness, as suggested in [53].
These results indicate that DetectCodeGPT is a model-free method that is robust against model discrepancies, making it a viable solution for real-world applications where the source model could be unknown or inaccessible.
6 Discussion
6.1 Why is DetectCodeGPT Effective?
@@
6.1.1 Preservation of Code Correctness
@@
6.1.2 Emulation of Human Randomness
As discussed in Section 3.4.5, human inherently exhibit less naturalness and more randomness in their use of stylistic tokens such as spaces and newlines than machines. For example, a human programmer may freely insert whitespace, especially newlines, in the code as they deem fit, whereas a machine programmer usually tries to stylize the code in a more standardized and modularized manner. Our proposed perturbation strategy mimics human’s free usage of spaces and newlines, thereby making the perturbation more “random” as is desired according to Section 3.
6.2 Strength of DetectCodeGPT
DetectCodeGPT presents a perturbation strategy that is not only efficient but also maintains the structural integrity of the code while closely mirroring human coding styles. We summarize the strengths of DetectCodeGPT as follows:
6.2.1 Enhanced Efficiency
Since it’s necessary to perturb each text for multiple times to ensure the effectiveness of detection [11, 45], as verified in Section 5.7, perturbing code using masked language modeling with other pretrained models will lead to huge computational cost. On the contrary, our method will save much more time and bring more efficiency. This translates to a significant conservation of computational resources, thus expediting the detection process and enhancing scalability. Mathematically, if $T_{\text{DetectGPT}}$ and $T_{\text{Ours}}$ denote the computational times of DetectGPT and our method respectively, we know that $T_{\text{Ours}}\ll T_{\text{DetectGPT}}$.
6.2.2 Superior Zero-Shot Performance
DetectCodeGPT distinguishes itself with a zero-shot learning capability, enabling it to detect machine-generated code without the necessity for training on extensive datasets. This model-agnostic advantage means that it can be generalized across various code LLMs and can still remain robust even when applied to unseen models, as demonstrated and verified in Section 5.8. In comparison to the supervised methods, which may suffer from data drift or the need for constant updates, our approach is not only more adaptable but also delivers better performance across various data sources. This makes DetectCodeGPT an agile and reliable tool for machine-generated code detection in a variety of contexts.
6.3 Limitations and Future Directions
@@
Looking ahead to future work, we plan to further improve the effectiveness of DetectCodeGPT when detecting machine-generated code at higher levels of generation randomness, such as $T$=$1.0$. We note from Table IV that although DetectCodeGPT outperforms other baselines at $T$=$1.0$, there is still large room for improvement. Although ensembling multiple detection models may help to improve the detection performance [53], we look forward to exploring more effective perturbation strategies based on code style to further enhance the detection efficacy.
@@
Additionally, the development of watermarking strategies [54] for machine-generated code represents an intriguing avenue for future research. Given that our method can effectively identify machine’s code, it could be adapted to assist in the creation of watermarking techniques. For instance, embedding a watermark could be achieved by inserting newlines and spaces in a code-specific pattern, which would increase detectablility without compromising functionality.
7 Related Work
@@
Recently, there has been much effort in detecting machine-generated text [9, 55]. The two main categories of detection methods are zero-shot and training-based methods, and our DetectCodeGPT falls into the former category, which eliminates the need for training data and brings more generalization ability.
@@
As for zero-shot methods, they are usually based on the discrepancy between likelihood and rank information of human and machine’s texts [39, 38, 40]. Leveraging the hypothesis in DetectGPT [11] that machine-generated text often has a negative curvature in the log probability when the text is perturbed, many perturbation based methods have been proposed [11, 45, 56]. These methods usually perturb the text by masking a continuous span of tokens and then recover the perturbed text using another LLM like T5 [29]. The benefit of these methods is that they are zero-shot and can be applied to any LLM without access to training data. However, the perturbation process is time-consuming and computationally expensive. When it comes to the training-based methods, fine-tuning the RoBERTa[49] or T5 [29] model with data collected from different model families at different decoding settings is a common practice [15, 57, 58]. Additional information like graph structure [59], perplexity from proxy models [60] have been shown to be helpful for detection. Moreover, techniques like adversarial training [61] and contrastive learning [59] have also been proposed to improve the detection performance. The main challenge of training-based methods is that they often lack generalization ability and require access to training data from the target model [11].
@@
However, research on identifying machine-generated code remains relatively scarce and is purported to be more challenging than discerning machine-produced text [13]. GPTSniffer was first proposed to detect machine-generated code with supervised CodeBERT training [47]. And the predominant strategies in current studies largely revolve around watermarking techniques, which embed unique markers into the code either during the training of the model or at the time of code generation [13, 54]. The detection of these watermarks subsequently enables the recognition of code generated by machines. It should be noted, however, that these watermarking methods are primarily designed to address issues related to code licensing and plagiarism [62, 63]. Their reliance on modifications to the generation model renders them unsuitable for general code detection tasks.
8 Conclusion
@@
In this paper, we have explored the nuanced differences between machine-generated and human-written code to enhance the trustworthiness of software development in the era of AI. By pioneering a comparative analysis across three aspects of code, namely, naturalness, style, and correctness, we have provided new insights into code provenance that machine tends to write more concise and natural code with a narrower spectrum of tokens, towards regularization to programming principles, and the disparity is more pronounced in stylized tokens such as whitespaces. Our proposed detection method, DetectCodeGPT, capitalizes on these insights, specifically targeting the stylized nuances of code with a novel perturbation strategy that is simple yet effective. The empirical success of DetectCodeGPT against several benchmarks confirms its potential to bolster security measures in programming and establish a more symbiotic relationship between human developers and LLMs. Our work marks a significant step towards responsible AI use in software engineering, ensuring clarity in code authorship and maintaining the integrity of code as a craft.